# Deep Learning Reduced Order Modelling on Parametric and Data driven domains

### v1

@@ Abstract
Partial differential equations (PDEs) are extensively utilized for modeling various physical phenomena. These equations often depend on certain parameters, necessitating either the identification of optimal parameters or solving the equation across multiple parameters to understand how a structure might react under different conditions. Performing an exhaustive search over the parameter space requires solving the PDE multiple times, which is generally impractical. To address this challenge, Reduced Order Models (ROMs) are constructed from a snapshot dataset comprising parameter-solution pairs. ROMs facilitate the rapid solving of PDEs for new parameters.
Recently, Deep Learning ROMs (DL-ROMs) have been introduced as a new method to obtain ROM. Additionally, the PDE of interest may depend on the domain, which can be characterized by parameters or measurements and may evolve with the system, requiring parametrization for ROM construction. In this paper, we develop a Deep-ROM capable of extracting and efficiently utilizing domain parametrization. Unlike traditional domain parametrization methods, our approach does not require user-defined control points and can effectively handle domains with varying numbers of components. Moreover, our model can derive meaningful parametrization even when a domain mesh is unavailable, a common scenario in biomedical applications. Our work leverages Deep Neural Networks to effectively reduce the dimensionality of the PDE and the domain characteristic function.
Keywords Reduced Order Modeling  $\cdot$ Deep Learning  $\cdot$ Partial Differential Equations
@@ 1 Introduction
Applications such as design optimization and real-world simulation often require solutions to partial differential equations across various parameters [1], [2], [3], [4]. Typically, with given parameters $\bm{\lambda}$ the objective is to find a function $\bm{u}$ such that:
$\mathcal{L}(\bm{u};\bm{\lambda})=f(\cdot;\bm{\lambda}),\text{ on $\Omega(\bm{\lambda})$}$ (1)
$\mathcal{B}(\bm{u};\bm{\lambda})=g(\cdot;\bm{\lambda}),\text{ on $\partial\Omega(\bm{\lambda})$},$ (2)
where $\mathcal{L}$ and $\mathcal{B}$ are differential operators, $f$ and $g$ are functions and $\Omega$ is the domain, and all of these can depend on the parameter vector $\bm{\lambda}$. This task can be formulated as finding a mapping from parameters to their corresponding solutions. A naive approach would be to just solve PDE for each parameter value we want to examine. This can often be prohibitive due to time and memory constraints. To mitigate this problem, a Reduced Order Model (ROM) is constructed, which relies on the assumption that all solutions for a specified parametric PDE belong to a lower-dimensional space. Classically, the problem of constructing ROM then involves determining the reduced basis [5]. In order to construct the ROM, one needs to obtain a dataset of high-fidelity parameter-solution pairs. This is usually done by employing the Galerkin finite element method (FEM) [6] which approximates solutions within a finite-dimensional space and uses a weak formulation. To solve a single instance of a PDE, the solution is approximated in a finite-dimensional function space. The problem of solving the PDE is then reformulated as the problem of finding the coefficients of the basis functions in that space. This results in a large system of equations and the discretized solution of PDE can be written as a large vector. To solve for many parameter values, one constructs and uses ROM. Its development consists of two phases: offline phase in which the ROM is constructed, and online phase, where the solution for the new parameter is evaluated. To construct ROM, one collects the dataset $(\bm{\lambda},\bm{u}_{i}),i=1,\ldots,N$, where $\bm{\lambda}$ is the i-th vector of parameters and $\bm{u}_{i}$ is the full-order model (FOM) solution corresponding to these parameters. The dataset comes from measurements or high-fidelity numerical solvers. Usually, using this data, linear reduction techniques are used to project solutions and equations into lower-dimensional space to obtain the ROM. To get a solution for the new parameter, it is then only necessary to solve in lower-dimensional space, which is faster than solving the original PDE. However, evaluation still requires solving a smaller system of equations. Furthermore, classical ROMs depend on the quality of the linear projection. This may result in a high number of equations for a reduced model. Because of that, few nonlinear methods were developed [7], [8], [9].
Regarding classical Reduced Order Models (ROMs), constructing ROMs that are effective across different domains presents several challenges. Firstly, to identify the projection onto the appropriate lower-dimensional space, all components of all solutions must have matching locations on the computational domain. This requirement is often unfeasible because irregular domains are discretized differently, making it impossible to ensure that all solutions have the same dimensionality. A common solution to this issue is to create a mapping between a fixed reference domain and the domain of interest, transforming the equation to apply to the fixed reference domain [10]. However, this approach is highly problem-dependent and not universally applicable. Additionally, domains may be derived from measurements, have varying numbers of subdomains or holes, or lack an available mesh. In such cases, it is unclear how to parametrize these domains and construct an appropriate mapping to a reference domain.
In this work, we address these challenges by employing convolutional autoencoder architecture on characteristic functions represented as a bitmap image to extract domain parametrization. This approach works in cases where the mesh is not available, when domains come from measurements, or when we have varying numbers of components or holes. The extracted parametrization is then incorporated into a Deep-ROM architecture. We demonstrate the effectiveness of our approach across various 2D problems.
@@ 2.1 Reduced Order Model via Proper orthogonal decomposition
Proper orthogonal decomposition (POD) [15] is a linear reduction technique. The basic assumption is that even though discretized solutions belong to a high-dimensional vector space of dimension $m$, solutions can be well approximated by a lower-dimensional vector space of dimension $k\ll n$. In this setting, all solutions can be vectorized and assembled into a matrix $\mathbf{U}\in\mathbf{M^{mn}}$. To obtain the basis of such a lower-dimensional vector space, SVD decomposition [16] is calculated, which means that orthogonal matrices $\mathbf{\hat{U}}\in\mathbf{M^{m}},\mathbf{\hat{V}}\in\mathbf{M^{n}}$ orthogonal matrices and diagonal $\Sigma\in\mathbf{M^{mn}}$ with non-negative decreasing numbers on its diagonal $\sigma_{1}\geq\sigma_{2}\geq\ldots\geq\sigma_{m}$ are found such that
@@
where $\bm{x}$ is $k$-dimensional vector. Now, instead of finding $\bm{u}$, we have to find $\bm{x}$. Then, when $Vx$ is substituted instead of $\bm{u}$ in the Galerkin FOM formulation, a new system with fewer unknowns is obtained. In Linear PDE, this results in decreasing the number of equations. In non-linear cases, this cannot be trivially done, so techniques such as the Discrete Empirical Interpolation Method (DEIM) [17] are developed. Further difficulties in employing this ROM arise when the domain depends on parameters because all discretized solutions need to have the same dimensionality and each coordinates of all solutions have to correspond to the same points in the domain. To obtain ROM on parametric domains $\Omega(\bm{\lambda})$, equations are mapped on reference domain $\hat{\Omega}$. Mapping $\mathcal{M}(\cdot;\bm{\lambda})$ maps the reference domain into $\Omega(\bm{\lambda})$, so the correspondence of points across different domains is established. To construct such a mapping, one can use free-form deformation (FFD) [18], [19], [20], radial basis function interpolation (RBF) [21], [22], inverse distance weighting interpolation (IDW) [23], [24], [25], or surface registration using currents [26]. Furthermore, domains should be parametrized as well. To get parametrization, a set of control points $\bm{x}_{1},\bm{x}_{2},\ldots,\bm{x}_{c}$ around the domain is selected, and displacements of these points from their corresponding positions on the reference domain are calculated. Control points on each domain have to match the same points in the reference domain through appropriate mappings to get meaningful parameters. Control points have to be chosen carefully so that they capture the shape characteristics well. This imposes difficulties when one wants to obtain one ROM for situations where there are a varying number of components or holes.
@@ 2.3 Deep ROM (DL-ROM)
N. Franco et al. introduced Deep ROM (DL-ROM) in [29]. The idea is to obtain ROM by exploiting autoencoder architecture [30]. The model consists of three neural networks: an encoder $\Psi_{E}^{S}$, a decoder $\Psi_{D}^{S}$, and a multi-layer perceptron (MLP) [31] $\Phi_{S}$. To obtain ROM, the encoder-decoder pair is trained to minimize reconstruction error:
@@
The encoder for standard DL-ROM $\Psi_{E}^{S}(\bm{u}_{i})$ maps the solution to a lower-dimensional space, called encoding, and the decoder lifts encoding to the original solution size. The autoencoder is then trained to minimize reconstruction error. Finally, $\Phi_{S}$ is trained to minimize the mean-squared error between its output and encoder encoding:
@@
Solution $\bm{u}$ for a new value of parameter $\bm{\lambda}$ is then:
@@
The whole architecture is shown on Figure 1
Figure 1: Basic DL-ROM architecture: encoder $\Psi_{E}^{S}$ and decoder $\Psi_{D}^{S}$ are trained to minimize reconstruction error between full order model solution and its reconstruction, After training, encodings of solutions are calculated with encoder $\Psi_{E}^{S}$ and Multilayer perceptron (MLP) $\Phi_{S}$ is trained to map parameters to encoding. To obtain a solution for new parameters $\bm{\lambda}$, parameters are passed through $\Phi_{S}$ and then through the decoder $\Psi_{D}^{S}$.
Notice that getting solutions from this kind of reduced-order model doesn’t require solving any additional equations and can be efficiently done for multiple parameters at the same time, suing GPU parallel computing. In this work, Convolutional Neural Networks (CNN) [32] were used for the encoder and decoder and all domains involved were rectangular.
@@ 2.4 Deep Learning Proper Orthogonal Decomposition ROM
Deep Learning Proper Orthogonal Decomposition ROM (POD-DL-ROM) was introduced in [33]. The procedure goes as follows. Solutions $\bm{u}_{i},i=1,\ldots N$ are reshaped into vectors of length $N_{h}$ and assembled into a matrix $\mathbf{A}$ with $N_{h}$ rows and $N$ columns. Columns of matrix $\mathbf{A}$ are $\bm{a}_{i},i=1,2,\ldots N$. The essence of the POD method is to get orthonormal vectors $\bm{\Phi_{i}},i=1,2,\ldots k$, $k\ll N$ vectors that maximize:
$\frac{1}{N}\sum_{i=1,\ldots N,j=1,\ldots k}|\langle\bm{a}_{i},\bm{\Phi}_{j}\ \rangle|^{2}$
So, the following approximation hold:
@@
Matrix $\mathbf{V}_{k}$ is usually found using SVD decomposition or randomized SVD (rSVD) [34]. So, the solution $u$ can be approximated with POD basis and represented as $k$-dimensional vector $[\langle\bm{u},\bm{\Phi}_{1}\rangle,\ldots\langle\bm{u},\bm{\Phi}_{k}\rangle]$. Then, similar to the black-box DL-ROM, the model consists of three networks: encoder $\Psi_{E}$, decoder $\Psi_{D}$, and parameter-encoding network $\Phi$. The difference is that the encoder and decoder are trained on POD coefficients, thus reducing the dimension of the problem. For this model to work on multiple domains, POD coefficients at the reference domain and mapping between domain and reference domain are needed.
@@ 2.5 Graph convolutional autoencoder approach for DL-ROMs
Following DL-ROM framework, in [35] Graph convolutional autoencoders (GNNs) [36], [37], [38], [39] are used for dimension reduction. Using the proposed autoencoder technique, one can obtain ROM for non-rectangle domains using the GNN autoencoder. In the proposed approach, graph pooling and a fully-connected layer work only when all domains have the same number of nodes and domain parametrization is known as well. In applications, especially in biomedicine, both constraints may not hold.
@@ 3 Our method
In this work, we develop an efficient Reduced Order Model (ROM) that operates effectively across different domains, based on a Deep Learning Reduced Order Model (DL-ROM) approach. We refer to our method as Deep Learning domain aware ROM (DL-DA ROM). To achieve this, meaningful parametrization of domains is essential. Domains can be described parametrically, may evolve as part of the system (e.g., moving domains), or can be obtained from measurements.
Since the ROM for given parameters has to predict the domain and values of solutions on that domain, we separate those two tasks. In this work, we assume that the solution is at least continuous and can be interpolated onto rectangular mesh without significant loss of accuracy, see Figure 2. By interpolating in this manner, Convolutional neural networks (CNNs) can be utilized to construct an autoencoder for solutions. This approach is particularly beneficial in scenarios where the domain is derived from measurements and meshing is prohibitively expensive.
Figure 2: Sample interpolated onto equidistant rectangular grid (on the left). Sample (on the right).
When domains are parametrically described, it is not necessary to assume that each domain has the same number of parameters; they may have varying numbers of subdomains or holes. Our approach assumes that each domain characteristic function can be interpolated onto a rectangular grid, see Figure 3.
Figure 3: Mesh representation of the domain (left), an image of domain characteristic function (right).
The domain of the problem can be derived either from measurements or parameters. In our proposed method, its characteristic function is described as a bitmap image, with pixels assigned a value of one or zero depending on whether the pixel is inside or outside the domain. A convolutional autoencoder is trained using binary cross-entropy loss [40] to classify each pixel as part of the domain or not which is given by
@@
where $\bm{y}$ is target bitmap image with $n$ rows and $m$ columns and $\bm{x}$ is network output image, whose pixels values $x_{ij}$ denote probability that pixel $x_{ij}$ is equal to one. The encoded domains are then used as the parametrization of the domain. The domain autoencoder is inspired by semantic segmentation encoder-decoder models in computer vision [41], [42], with two classes indicating whether a pixel is inside or outside the domain. This approach can be extended to problems with multiple domains, such as fluid-structure interaction problems, providing simultaneous geometric parametrization for the entire system. If domains can be generated synthetically, it is possible to use more samples for the domain autoencoder than for the solution autoencoder.
Domain autoencoder is shown in Figure 4.
Figure 4: Domain Autoencoder used for domain parametrization. Input to the convolutional neural network is a characteristic function represented as a bitmap image. The output of the autoencoder is a rectangular grid with the probability that a given pixel belongs to the domain. The encoder output is used as domain parametrization.
If the domain is either known from measurements or is parametrized, the encoding of the domain can be used as additional parameters in the network $\Phi$. The model is shown in Figure 5 and all important terms in this work are given in table 1.
[sca]
Figure 5: Deep ROM for parametric domains: In addition to vanilla Deep ROM, we introduce domain autoencoder, which extracts parameters for the domain. These encodings are then used as additional parameters in Deep ROM.
label meaning $\Omega(\bm{\lambda})$ domain of the problem, in our work, is dependent on parameters $\Psi_{E}^{S}$ solution encoder $\Psi_{D}^{S}$ solution decoder $\Phi_{S}$ neural network that maps parameters into solution encoding $\Psi_{E}^{D}$ domain encoder $\Psi_{D}^{D}$ domain decoder
@@
Constructing the ROM, also known as the offline phase, involves gathering the data, either with numerical simulation, measurements, or possibly a combination of both. Domain and solution autoencoders are trained to obtain meaningful low-dimensional encodings. Before training the solution autoencoder training solution data is standardized by applying the following transformation:
$\bm{u}_{i}=\frac{\bm{u}_{i}-\mu_{s}}{\sigma_{s}},1,\ldots N$
where $N$ is a number of samples in the training dataset, $\mu_{s}$ and $\sigma_{s}$ are mean and standard deviation of the training dataset After training the solution and domain autoencoders, domain encoding, and other equation parameters are concatenated. Training parameters are standardized by applying the following transformation:
$\lambda_{ij}=\frac{\lambda{ij}-\mu_{pj}}{\sigma_{pj}},i=1,\ldots N,j=1\ldots k$
where $k$ is dimension of parameter vector, $N$ number of samples and $\mu_{pj},\sigma_{pj}$ are mean and standard deviation of each parameter separately. Afterward, the network $\Phi_{S}$ is trained to predict solution encodings from parameters.
The procedure is summarized in Algorithm 1. After all data collection and training are completed, solutions for unseen parameters are calculated in the online phase, as detailed in Algorithm 2. For new values of equation parameters and domain, this involves passing the domain through the domain encoder $\Psi_{E}^{D}$ concatenating the resulting encoding with the equation parameters, passing them through the MLP $\Phi_{S}$ and finally through the solution decoder $\Psi_{D}^{S}$ to obtain prediction $\hat{u}$. The final ROM solution is then given by:
$\bm{u}=\sigma_{s}\bm{u}+\mu_{s}$
@@
Save mean $\mu_{s}$ and standard deviation $\sigma_{s}$ of training dataset and mean $\mu_{p}$ and standard deviation $\sigma_{p}$ of training feature dataset

### v2

# Reduced Order Modeling of Partial Differential Equations on Parameter-Dependent Domains Using Deep Neural Networks

@@ Abstract
Partial differential equations (PDEs) are widely used for modeling various physical phenomena. These equations often depend on certain parameters, necessitating either the identification of optimal parameters or the solution of the equations over multiple parameters. Performing an exhaustive search over the parameter space requires solving the PDE multiple times, which is generally impractical. To address this challenge, reduced order models (ROMs) are built using a set of precomputed solutions (snapshots) corresponding to different parameter values. These models allow for a fast approximation of the PDE solution when new parameters are introduced.
Recently, Deep Learning ROMs (DL-ROMs) have been introduced as a new method to obtain ROM, offering improved flexibility and performance. In many cases, the domain on which the PDE is defined also varies and may be described by parameters or measurements. Capturing this variation is important for building accurate ROMs but is often difficult, especially when the domain has a complex structure or changes topology.
In this paper, we propose a Deep-ROM framework that can automatically extract useful domain parametrization and incorporate it into the model. Unlike traditional domain parameterization methods, our approach does not require user-defined control points and can effectively handle domains with varying numbers of components. It can also learn from domain data even when no mesh is available, which is often the case in biomedical applications. Using deep neural networks, our approach reduces the dimensionality of both the PDE solution and the domain representation, making it possible to approximate solutions efficiently across a wide range of domain shapes and parameter values. We demonstrate that our approach produces parametrizations that yield solution accuracy comparable to models using exact parameters. Importantly, our model remains stable under moderate geometric variations in the domain, such as boundary deformations and noise—scenarios where traditional ROMs often require remeshing or manual adjustment.
Keywords Reduced Order Modeling $\cdot$ Deep Learning $\cdot$ Partial Differential Equations $\cdot$ Parameter-Dependent Domains
@@ 1 Introduction
Applications such as design optimization and simulations in real-world settings often require solving partial differential equations (PDEs) for a wide range of parameter values [1], [2], [3], [4]. Typically, given a parameter vector $\bm{\lambda}$, the goal is to find a solution $\bm{u}$ such that:
$\mathcal{L}(\bm{u};\bm{\lambda})=f(\cdot;\bm{\lambda})\text{ on $\Omega(\bm{\lambda})$},$ (1)
$\mathcal{B}(\bm{u};\bm{\lambda})=g(\cdot;\bm{\lambda})\text{ on $\partial\Omega(\bm{\lambda})$},$ (2)
where $\mathcal{L}$ is a differential operator, $\mathcal{B}$ is a boundary operator, $f$ and $g$ are the source and boundary data, and $\Omega$ is a parameter-dependent domain. This task can be viewed as learning a solution map from the parameters $\bm{\lambda}$ to their corresponding solutions $\bm{u}$. A straightforward approach is to numerically solve the PDE for each parameter instance, but this becomes computationally prohibitive as the number of evaluations increases.
To address this, reduced order models (ROMs) are commonly used to approximate the solution map with a much lower computational cost. These models are built on the assumption that solutions to the parametric PDE lie on a low-dimensional manifold [5]. The construction of such a ROM typically involves two phases. In the offline phase, a dataset of high-fidelity parameter–solution pairs $(\bm{\lambda}_{i},\bm{u}_{i})$ is collected using methods such as the finite element method (FEM) [6] and ROM is obtained. In the online phase, the ROM is evaluated on unseen parameters. Full-order solutions are then projected on a reduced basis, often obtained using techniques such as Proper Orthogonal Decomposition (POD) [7]. In the online phase, a reduced system is solved to obtain an approximate solution for a new parameter value, significantly reducing computation time.
However, standard ROMs face limitations, especially when applied to PDEs defined on parameter-dependent domains. Constructing a projection-based ROM typically requires all solutions to be discretized over a common mesh or aligned computational domain, so that basis functions are defined consistently across parameter instances. This assumption breaks down when the domain itself varies with the parameters, which is common in many real-world scenarios.
A typical workaround is to map each domain to a fixed reference configuration and transform the PDE accordingly [8]. However, this mapping is often problem-specific and not easily generalized. Although this is feasible when domains vary smoothly and share the same topology, it becomes problematic when domains differ structurally, for instance, when holes appear or disappear, or when the number of connected components changes. In such cases, defining a single reference domain and constructing consistent mappings becomes difficult, or even impossible.
To address these challenges, we propose a data-driven framework based on convolutional autoencoders to learn a compact domain representation directly from its characteristic function. We represent each domain as a binary image and use a convolutional autoencoder to extract a low-dimensional encoding. This encoding serves as a learned domain parametrization, which we then integrate into a neural reduced order model that predicts PDE solutions. This approach requires no meshing, no manual alignment, and no user-defined control points. It is naturally suited to problems where the domain is defined through measurements or where topological complexity varies across parameter instances. Moreover, in many real-world scenarios—such as biomedical simulations, flow over irregular surfaces, or domains derived from imaging—constructing a computational mesh is often impractical or infeasible. This limits the applicability of traditional ROM techniques that rely on consistent meshing, graph connectivity, or control-point mappings to handle domain variability.
This challenge is particularly acute when domains are derived from imaging or experimental data, where mesh or graph structures may not be available at all. To overcome these limitations, we propose a fully data-driven, mesh-free approach that operates directly on binary images representing domain characteristic functions. Using convolutional autoencoders, our method learns compact geometric encodings without requiring explicit parametrization or meshing.
We demonstrate the effectiveness of this approach on synthetic two-dimensional PDE problems designed to reflect practical challenges, including geometric variability, boundary noise, and structural differences across domains.
@@ 2.1 Reduced Order Model via Proper orthogonal decomposition
Proper orthogonal decomposition (POD) [7, 12] is a linear reduction technique. The basic assumption is that even though discretized solutions belong to a high-dimensional vector space of dimension $M$, solutions can be well approximated by a lower-dimensional vector space of dimension $k\ll M$. In this setting, all solutions can be vectorized and assembled into a matrix $\mathbf{U}\in\mathbf{\mathbb{R}^{M\times N}}$. To obtain the basis of such a lower-dimensional vector space, SVD decomposition [13] is calculated, which means that orthogonal matrices $\mathbf{\hat{U}}\in\mathbf{\mathbb{R}^{M\times M}},\mathbf{\hat{V}}\in\mathbf{\mathbb{R}^{N\times N}}$ and diagonal $\Sigma\in\mathbf{\mathbb{R}^{M\times N}}$ with non-negative diagonal elements $\sigma_{1}\geq\sigma_{2}\geq\ldots\geq\sigma_{min\{M,N\}}$, are found such that
@@
where $\bm{x}$ is $k$-dimensional vector. Now, instead of finding $\mathbf{u}$, we have to find $\mathbf{x}$. Then, when $\mathbf{Vx}$ is substituted instead of $\mathbf{u}$ in the Galerkin FEM formulation, a new system with fewer unknowns is obtained. In linear PDE, this results in decreasing the number of equations. In non-linear cases, this cannot be trivially done, so techniques such as the Discrete Empirical Interpolation Method (DEIM) [14] are developed. Further difficulties in employing this ROM arise when the domain depends on parameters because all discretized solutions need to have the same dimension and each coordinates of all solutions have to correspond to the same points in the domain. To obtain ROM on parametric domains $\Omega(\bm{\lambda})$, equations are mapped to the reference domain $\hat{\Omega}$. Mapping $\mathcal{M}(\cdot;\bm{\lambda})$ maps the reference domain into $\Omega(\bm{\lambda})$, so the correspondence of points across different domains is established. To construct such a mapping, one can use free-form deformation (FFD) [15], [16], [17], radial basis function interpolation (RBF) [18], [19], inverse distance weighting interpolation (IDW) [20], [21], [22], or surface registration using currents [23]. Furthermore, domains should be parametrized as well. To get parametrization, a set of control points $\bm{x}_{1},\bm{x}_{2},\ldots,\bm{x}_{c}$ around the domain is selected, and displacements of these points from their corresponding positions on the reference domain are calculated. Control points on each domain have to match the same points in the reference domain through appropriate mappings to get meaningful parameters. Control points have to be chosen carefully so that they capture the shape characteristics well. This imposes difficulties when one wants to obtain a single ROM for situations where there are a varying number of components or holes.
@@ 2.3 Deep ROM (DL-ROM)
Franco et al. introduced Deep ROM (DL-ROM) in [26]. The idea is to obtain ROM by exploiting autoencoder architecture [27]. The model consists of three neural networks: an encoder $\Psi_{E}^{S}$, a decoder $\Psi_{D}^{S}$, and a multi-layer perceptron (MLP) [28] $\Phi_{S}$. To obtain ROM, the encoder-decoder pair (i.e. autoencoder) is trained to minimize reconstruction error:
@@
The encoder for standard DL-ROM $\Psi_{E}^{S}(\bm{u}_{i})$ maps the solution to a lower-dimensional space, called encoding, and the decoder lifts the encoding to the original solution size. Finally, $\Phi_{S}$ is trained to minimize the mean square error between its output and encoder encoding:
@@
The solution $\bm{u}$ for a new value of parameter $\bm{\lambda}$ is then:
@@
The entire architecture is shown on Figure 1
Figure 1: Basic DL-ROM architecture: encoder $\Psi_{E}^{S}$ and decoder $\Psi_{D}^{S}$ are trained to minimize reconstruction error between full order model solution and its reconstruction, After training, encodings of solutions are calculated with encoder $\Psi_{E}^{S}$ and Multilayer perceptron (MLP) $\Phi_{S}$ is trained to map parameters to encoding. To obtain a solution for a new parameter $\bm{\lambda}$, parameters are passed through $\Phi_{S}$ and then through the decoder $\Psi_{D}^{S}$.
Notice that getting solutions from this kind of reduced-order model does not require solving any additional equations and can be efficiently done for multiple parameters at the same time, using GPU parallel computing. Authors used Convolutional Neural Networks (CNN) [29] for the encoder and decoder and all domains involved were rectangular.
@@ 2.4 Deep Learning Proper Orthogonal Decomposition ROM
Deep Learning Proper Orthogonal Decomposition ROM (POD-DL-ROM) was introduced in [30]. The procedure goes as follows. Solutions $\bm{u}_{i},i=1,\ldots N$ are reshaped into vectors of length $N_{h}$ and assembled into a matrix $\mathbf{A}$ with $N_{h}$ rows and $N$ columns. Columns of matrix $\mathbf{A}$ are $\bm{a}_{i},i=1,2,\ldots N$. The essence of the POD method is to get orthonormal vectors $\bm{\Phi_{i}},i=1,2,\ldots k$, $k\ll N$ that maximize:
$\frac{1}{N}\sum_{i=1,\ldots N,j=1,\ldots k}|\langle\bm{a}_{i},\bm{\Phi}_{j}\ \rangle|^{2}.$
Then, the following approximation hold:
@@
Matrix $\mathbf{V}_{k}$ is usually found using SVD decomposition or randomized SVD (rSVD) [31]. The solution $u$ can be approximated with POD basis and represented as $k$-dimensional vector $\langle\bm{u},\bm{\Phi}_{1}\rangle,\ldots\langle\bm{u},\bm{\Phi}_{k}\rangle]$. Then, similar to the black-box DL-ROM, the model consists of three networks: encoder $\Psi_{E}$, decoder $\Psi_{D}$, and parameter-encoding network $\Phi$. The difference is that the encoder and decoder are trained on POD coefficients, thus reducing the dimension of the problem. For this model to work on multiple domains, POD coefficients at the reference domain and mapping between domain and reference domain are needed.
@@ 2.5 Graph convolutional autoencoder approach for DL-ROMs
Following DL-ROM framework, in [32] Graph convolutional autoencoders (GNNs) [33], [34], [35], [36] are used for dimension reduction. Using the proposed autoencoder technique, one can obtain ROM for non-rectangle domains using the GNN autoencoder. However, graph pooling and a fully-connected layer work only when all domains have the same number of nodes and domain parametrization is known as well. In applications, especially in biomedicine, both constraints may not hold.
In [37], a harmonic mapping is used to map the domain and solution to a reference configuration. The neural operator is learned on a reference configuration and mapped to the original domain via inverse of the harmonic mapping. This approach may not be feasible when such harmonic mapping is difficult or impossible to construct due to a wide geometric or topological variability.
Similarly as in our work, in [38] a characteristic function of the domain is used, and the solutions are interpolated onto a rectangular grid using CNN. The input is a domain characteristic function and the output is the solution on that domain. Contrary to our work, this model is not capable of handling non-geometric parameters simultaneously.
@@ 3 Our method
In this work, we develop an efficient Reduced Order Model (ROM) that operates effectively across different domains, based on a Deep Learning Reduced Order Model (DL-ROM) approach. We refer to our method as Deep Learning Domain Aware ROM (DL-DA-ROM). A key part of this approach is learning a meaningful parametrization of the domain. In many problems, the domain may be defined through parameters, evolve with the system (e.g., moving domains), or come from measurements.
Since the ROM needs to predict both the domain and the solution on that domain, we separate those two tasks. We assume that the solution is at least continuous and can be interpolated onto rectangular grid without significant loss of accuracy, see Figure 2. This interpolation allows us to use CNNs to build an autoencoder for the solutions. This setup is especially useful when domains come from measurement data and creating a mesh is either difficult or too costly. While it is possible to interpolate the solution and then mask out the region outside the domain, we found that using the full interpolated solution yields better performance. Interpolation naturally smooths the data, reducing sharp discontinuities introduced by irregular meshes or complex geometries. This smoothing effect makes the solution easier to approximate with CNNs, which are particularly well-suited for capturing smooth spatial patterns. As a result, CNN-based autoencoders can more efficiently learn a compact representation of the solution field without being disrupted by artificial discontinuities at domain boundaries.
Figure 2: Left panel shows a sample function, and the right panel shows its interpolation into an equidistant rectangular grid.
When domains are parametrically described, it is not necessary to assume that each domain has the same number of parameters; they may have varying numbers of subdomains or holes. Our approach assumes that each domain characteristic function can be interpolated onto a fixed rectangular grid, see Figure 3.
Figure 3: Comparison between a mesh-based domain representation (left) and the corresponding interpolated characteristic function represented as an image used by our method (right). The binary image representation allows us to learn a low-dimensional domain encoding without requiring a mesh.
To represent the domain, we use its characteristic function interpolated onto a rectangular grid, resulting in a binary image. This image is stored as a bitmap, where each pixel is assigned a value of 1 if the corresponding point lies inside the domain and 0 otherwise. A convolutional autoencoder is then trained on these bitmap images to extract a compact latent representation, which serves as a learned, low-dimensional domain parameterization. This mesh-free approach accommodates complex geometries and topological variations without requiring control points, meshing, or predefined mappings. To extract the domain parametrization, we train a convolutional autoencoder on these binary images, using the binary cross-entropy loss[39] to classify each pixel as part of the domain or not which is given by
@@
where $\bm{y}$ is target bitmap image with $n$ rows and $m$ columns and $\bm{x}$ is network output image, whose pixels values $x_{ij}$ denote probability that pixel $x_{ij}$ is equal to one. The encoded domains are then used as the parametrization of the domain. The low-dimensional latent code produced by the encoder is used as domain parametrization in our model. The autoencoder architecture is inspired by the encoder-decoder networks commonly used in semantic segmentation [40, 41], with two classes that distinguish whether each pixel lies within or outside the domain. This approach can be extended to problems with multiple domains, such as fluid-structure interaction problems, providing simultaneous geometric parametrization for the entire system.
If domains are synthetically generated, it is also possible to train the domain autoencoder with a larger dataset than the solution autoencoder, which can help improve the quality of the learned parametrization and its generalization to unseen geometries.
The domain autoencoder is shown in Figure 4.
Figure 4: Domain Autoencoder used for domain parametrization. The input to the convolutional neural network is a characteristic function represented as a bitmap image. The output of the autoencoder is a rectangular grid where each pixel value represents the predicted probability that the pixel belongs to the domain. The low-dimensional encoder output (latent representation) is used as the domain parametrization.
Fully-connected neural network $\Phi_{S}$ is trained to map input parameters (e.g., domain encoding and physical parameters) to the solution representation. The predicted latent solution is then passed through the decoder $\Psi^{S}_{D}$ to reconstruct the full solution on the interpolated grid. The model is shown in Figure 5 and all important terms are given in Table 1.
Figure 5: Deep ROM architecture for PDEs with parameter-dependent domains. In addition to the standard Deep ROM structure, our method incorporates a domain autoencoder (bottom) that encodes the characteristic function of the domain into a low-dimensional latent representation. This representation is combined with other problem parameters and passed to a fully connected network $\Phi_{S}$, which predicts the solution encoding. The predicted latent solution is then decoded using $\Psi^{S}_{D}$ to reconstruct the solution on the full-order model (FOM) grid. Both the solution and domain autoencoders are trained using CNNs. The domain representation serves as a learned geometric parametrization that allows the ROM to generalize across different domain shapes and configurations.
label meaning $\Omega(\lambda)$ parameter-dependent domain $\Psi_{E}^{S}$ solution encoder $\Psi_{D}^{S}$ solution decoder $\Phi_{S}$ neural network that maps parameters into solution encoding $\Psi_{E}^{D}$ domain encoder $\Psi_{D}^{D}$ domain decoder
@@
Constructing the ROM-also known as the offline phase, involves gathering the data, either using numerical simulation, measurements, or possibly a combination of both. During this phase, both the domain and solution autoencoders are trained to obtain meaningful low-dimensional representations. Before training the solution autoencoder training solution data is standardized by applying the following transformation:
$\bm{u}_{i}=\frac{\bm{u}_{i}-\mu_{s}}{\sigma_{s}},1,\ldots N,$
where $N$ is the number of samples in the training dataset and $\mu_{s}$ and $\sigma_{s}$ are the mean and standard deviation of the training dataset. After the solution and the domain autoencoders are trained, we evaluate the domain encoder to extract the low-dimensional domain parameters. These are then concatenated with the PDE parameters. Prior to training the solution prediction network $\Phi_{S}$, the concatenated input equation and domain parameters are standardized:
$f_{ij}=\frac{f_{ij}-\mu_{j}^{p}}{\sigma_{j}^{p}},i=1,\ldots N,j=1\ldots k,$
where $f_{i,j}$ the $j$-th component of the concatenated parameter vector for the $i$-th training sample, $k$ is dimension of a parameter vector, $N$ number of samples and $\mu_{j}^{p},\sigma_{j}^{p}$ are mean and standard deviation of the $j$-th component of parameter vectors in the training set. The network $\Phi_{S}$ is then trained to map from this combined, standardized parameter vector to the corresponding solution encoding. The full procedure is summarized in Algorithm 1. Once the training is complete, solutions for unseen parameters are calculated in the online phase, as detailed in Algorithm 2. Given a new domain and PDE parameters, the domain is first passed through the trained encoder $\Psi_{E}^{D}$, the standardized result is concatenated with the standardized PDE parameters, and the solution is reconstructed using the trained networks. The predicted ROM solution $\hat{u}$ is then transformed back to the original scale using:
$\bm{\hat{u}}=\sigma_{s}\bm{\hat{u}}+\mu_{s}.$
@@
Save mean $\bm{\mu}_{s}$ and standard deviation $\bm{\sigma_{s}}$ of training dataset and mean $\bm{\mu}_{p}$ and standard deviation $\bm{\sigma}_{p}$ vectors of training feature dataset